# The building blocks of software work explain coding careers and language popularity

### v1

@@ 2 Results
2.1 Constructing a space of software tasks
Software development tasks can be viewed as recurring categories of programming problems developers encounter. SO, with its vast repository of user-generated questions, offers a unique window into these real-world challenges. To organize this repository and facilitate retrieval, each question is labeled with one or more of 66,000 curated tags that indicate the specific tools, frameworks, or concepts involved.
We use these tag–question relationships to group related issues into “canonical” software tasks, applying a bipartite stochastic block model (SBM) (Peixoto, 2017, 2014; Gerlach et al., 2018) to detect communities in the graph of questions and tags (Fig. 1a). We ignore tags that refer to programming languages and use them later on to study how tasks are distributed across languages. To generate concise names for each task community, we prompt ChatGPT-4.0 for labels that capture the essence of the tags comprising each group (see Appendix B). This process yields a set of 237 software tasks, such as Develop NLP models and tools for text analysis, Establish secure server access and control, and Deploy and secure a scalable web app.
Figure 1: Mapping software tasks. a. Stylized depiction of the bipartite question-tag network. SBM groups tags into communities (tasks) that connect to similar sets of questions. ChatGPT-4.0 finds a common label that summarizes each community’s tag information. b. Task space. Pointwise mutual information ($\mathrm{PMI}$) expresses how surprisingly often two tasks are performed by the same users. UMAP embeds the resulting co-occurrence network in a 2-dimensional plane (the task space). c. Close-up on Develop AI models task, depicting the original network structure among the 10 most closely related tasks. d. Task values. Nodes are colored according to their task value, estimated from salary information in the SO 2023 developers survey. Darker shades indicate more valuable tasks. e. Table of the five most and least valuable tasks. For details, see Methods.
Because certain tasks complement each other, they often co-occur within the same job roles (Anderson, 2017; Alabdulkareem et al., 2018; Neffke, 2019; Stephany and Teutloff, 2024). To investigate this task bundling, we identify which tasks are frequently performed by the same SO users. This yields a network that links tasks based on their co-occurrence in user activity. To reduce dimensionality and facilitate visualization, we then embed this network in a two-dimensional plane.
The resulting task space contains several clusters of related tasks, each shown in a different color in Fig. 1b (see Methods). Many of these clusters group tasks with comparable labor market returns. This is illustrated in Fig. 1d, where, we plot task values inferred from self-reported wages in a large-scale survey run among SO users (see Methods). On the right side of the task space, for example, three clusters revolve around website development, typically linked to lower wages. By contrast, tasks in the top center—focusing on advanced iOS and Android development—offer significantly higher earnings. Additional high-value clusters include those for AI and machine learning (AI/ML), Advanced Programming Concepts, Cloud Computing, and DevOps. The table of Fig. 1e lists the five highest-paying tasks, three of which are related to iOS app development, and five lowest-paying tasks, all associated with more basic web development.
Notably, clusters that are conceptually linked—such as AI/ML and Statistics and Data Analysis, and iOS and Android development—tend to lie near each other in the task space. To investigate the network structure of the task space, we can examine individual tasks in greater detail. For example, focusing on the ten tasks most closely related to developing AI models, Fig. 1c reveals strong connections to other statistical tasks, as well as tasks involving image and natural language processing, data visualization, and packaging models into standalone applications.
2.2 Validation using job ads
How well do these software tasks capture real-world development work beyond SO? To find out, we examine approximately 50,000 software job ads from HackerNews [^1] https://www.hackernews.com/ (see Fig. 2a and Methods). Using an LLM, we parse each job ad’s text to extract salaries and the list of tasks required of successful candidates.
Next, we convert each job’s extracted tasks into a 237-dimensional binary vector, where each dimension corresponds to one of the SO tasks identified previously. We accomplish this by embedding both the job’s task descriptions and the SO task labels into a 384-dimensional space using sBERT(Reimers and Gurevych, 2019). We then match each extracted task to its closest SO task by comparing cosine similarities. In this way, each job ad becomes a vector of job requirements, expressed in terms of SO tasks. Finally, we combine this vector with task-value information and measures of task relatedness. Figure 2a (see Methods) illustrates this process.
Figure 2: Job ads. a. Schematic representation of the workflow to extract salary and task requirements from online job ads by prompting ChatGPT. Task requirements are converted to the 237-dimensional SO task vectors of Fig. 1 based on cosine similarities between text embeddings of task requirements and SO tasks. b. Prediction of task requirements. 40% of task vector elements are masked and grouped into 10 equally sized bins based on the fit of the masked task to the (unmasked) task requirements of a job (see Methods). The plot displays the estimated probability that tasks in a bin are required in the job. c. Predictions of wage offers. Jobs are grouped into equally sized bins based on the average value of required tasks. The vertical axis shows the estimated mean of the advertised wage offers. Vertical bars in panels b and c represent 95% confidence intervals.
The task space strongly predicts which tasks appear together in job ads. To demonstrate this, we combine all job-task vectors into a single dataset and randomly mask 40% of the entries. For each masked task, we then measure its relatedness to the remaining unmasked tasks for that job. As shown in Fig. 2b, the probability that a masked task is indeed required (i.e., has a “1” in the task vector) increases sharply with its relatedness to the other tasks, rising 70-fold from below 1% for unrelated tasks to over 12% for strongly related ones.
@@ 2.2 Validation using job ads
The tasks we identify also predict wage offers (Fig. 2c). Salaries in job ads increase alongside the average value of the tasks a job requires, and regression analysis confirms this: a 10% rise in the average value of these tasks is associated with a 9% increase in the offered wage (see Appendix C.2).
2.3 Task dynamics: growth, diversification, feedback
Figure 3: Task dynamics. a. Task user-share change from 2009 to 2022. Purple markers signal increases, orange markers decreases, in user-shares between 2009 and 2022. Marker transparency reflects the size of shifts: darker tones indicate larger changes in user shares. b. Estimated probability of diversifying into new tasks at different values of density, users’ relatedness-weighted experience in other tasks: $d_{\theta,u}=\sum_{\kappa}\frac{\mathrm{R}_{\theta,\kappa}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}X_{\kappa,u}$, where $X_{\kappa,u}$ denotes user $u$’s experience in task $\kappa$ and $\mathrm{R}_{\theta,\kappa}$ the relatedness between tasks $\theta$ and $\kappa$. c. Regression analysis of answer popularity on task experience. Popularity is measured as the number of votes the answer receives, $\log(\#\text{votes}_{a}+1)$ or whether or not the answer is the top answer to the question, task experience as the number of prior answers the user provided to questions on task $\theta$ in the preceding two years, $\log(\#\text{answers}_{u(a),\theta,t(a)})$. Control variables include $\log(\#\text{answers}_{q(a)})$ and $\log\left(\sum_{\alpha\in A_{q}}\#\text{votes}_{\alpha}+1\right)$, the total number of answers provided to question $q(a))$, and the sum of all votes across these answers. To avoid problems due to $\log 0$ values, we add 1 to counts that can evaluate to 0. The plot shows point estimates with their 95% confidence intervals. Purple markers refer regression analyses of whether or not an answer is the top answer, orange markers to analyses of the number of votes. d. Regression of whether or not a user will adopt a new task $\theta$, on the task value, $V_{\theta}$, controlling for the user’s density around the task, $d_{\theta,u}$. Purple markers refer to Python-related questions only, orange markers to the full sample.
Figure 3a illustrates how SO activity shifted across software tasks between 2009 and 2022. Notable growth appears in Web Frameworks, certain Android-related tasks, Cloud Computing, and AI/ML. Conversely, tasks associated with Advanced Programming Concepts (e.g., Develop a data structure library for efficient data manipulation, Develop a multimedia application in Windows, or Develop a thread-safe application to prevent race conditions), as well as Networking, Enterprise and Web App Development, have seen their relative importance on SO decline.
We can also apply task vectors to study changes at the individual level, studying how individual SO users develop expertise over time. To show this, we generate a task vector for each user by counting how many answers they contribute in a two-year interval for each of the 237 defined tasks (see Methods). This approach captures the intensity of engagement in different areas of software work. We use these user task vectors to examine how developers add new skills to their profiles over time. In particular, we test whether task relatedness—reflected in the task space—predicts which new tasks a user will adopt on SO. As shown in Fig. 3b, users are far more likely to pick up tasks closely related to those they already know, with the probability of transitioning to a “new” task rising from 0.9% (when the task is unrelated) to 14.3% (when it is highly related).
We also expect that a user’s task-specific experience should correlate with the popularity of their answers. Testing this hypothesis is complicated, since users often decide to answer a question only after gauging existing solutions—a process that introduces selection bias. In Appendix D.1, we address these biases using instrumental variables. Here, we mitigate them partially by focusing on only the earliest answer posted to each question. We then ask whether users with greater experience in tasks connected to the question receive more votes. Concretely, we regress a binary outcome—whether the answer is the highest-voted—on the user’s prior task experience, controlling for the question’s total number of answers and sum of votes. As a robustness check, we also predict the total votes an answer receives.
Figure 3c reveals a statistically significant, positive link between users’ task experience and the popularity of their answers. The estimates indicate that a 10% rise in task experience boosts the likelihood of an answer becoming the top-voted solution by 0.1 pp (±$\pm 0.00004$) and increases the number of votes it receives by 0.14% (±$\pm 0.006\%$). In the SI, section D.2 we show that when with sample-selection corrections, the estimated effects grow to 0.006 pp (±$\pm 0.0035$) and 0.9% (±$\pm 0.41\%$), respectively.
Users also need motivation to adopt new tasks, whether for higher potential rewards or because the tasks align closely with their existing skills. To examine these motivations, we regress task entry on the estimated task value (from the aforementioned survey) and the user’s density around the task, i.e., the relatedness-weighted average experience in other tasks the user engages with (see Methods). As shown in Fig. 3d, developers do not universally pursue high-value tasks—likely due to the higher technological barriers or learning costs associated with them—and tend to choose tasks that are more closely related to their current skill sets (see Appendix D.3). However, Python users follow a different pattern: task value has a positive effect on their decision to adopt a new skill, suggesting that Python’s versatility allows them to more readily acquire profitable tasks. This highlights Python’s distinct role in the software ecosystem: as a general-purpose language, it enables users to branch into new, higher-value tasks more easily, altering and expanding their career trajectories.
2.4 Tasks and programming languages
@@
Programmers can draw on a wide variety of programming languages, each with different levels of suitability and popularity for specific software tasks. Our observation that Python users often diversify into higher-value tasks suggests that language choice significantly influences developers’ career paths. Yet debates persist about which languages best serve which tasks. Our task constructs promise a novel way to study language usage. To explore this, we use tags that refer to specific programming languages to associate SO posts with the languages they refer to (see Methods). This allows us to track which tasks a language supports.
@@
Figure 4a presents a matrix illustrating which programming languages (used by at least 10 SO users) are used in each of the 237 SO tasks. The matrix’s triangular shape signals task nestedness, with an NODF value of 81.5 (Payrató-Borràs et al., 2020), indicating that tasks can be roughly ranked from general to specialized based on how widespread they are across languages. Moreover, although different languages may be developed with specific use cases in mind, in practice, this has not led to a strict specialization where tasks are divided between dedicated programming languages. Instead, the most specialized tasks are typically handled by languages that cover almost the full spectrum of tasks, while languages used for only a few tasks are largely confined to the most ubiquitous ones.
@@
Figure 4: Programming languages. a. Task-language matrix. Elements are colored when at least 10 SO users have at least one answer post in the task-language combination. b. Number of tasks in which a programming language ranks as the top language in terms of SO users. The graph shows time-series for the largest six languages in terms of cumulative SO posts between August 2008 and June 2023. c. Python’s task footprint. Nodes are colored if Python ranks among the top 3 programming languages for the associated task. Fully colored: 1st rank, 50% transparency: 2nd rank, 75% transparency: 3rd rank.
@@
Figure 4b illustrates how different programming languages compete for tasks over time. Focusing on the six largest languages on Stack Overflow, it tracks how many tasks each language dominates by user counts. The most prominent trend is Python’s steady climb: from just nine tasks in 2009, it rises to become the top language for over 80 of the 237 tasks by 2022, surpassing well-established options like C# and JavaScript. In the SI, section F, we further demonstrate how languages compete over tasks by focusing on the abrupt shift in iOS programming from Objective-C to Swift.
@@
Figure 4c visualizes Python’s growth, highlighting all tasks in which it ranks among the top three languages. Initially focused on data science, Python spread between 2014 and 2018 into tasks ranging from app development to web-related work, DevOps, and cloud computing. Today, it is the dominant language for at least one task in every cluster. As detailed in Appendix E, this pattern is preserved when we re-weight languages by their GitHub script counts, a proxy for real-world usage—indicating that Python’s reach extends well beyond Stack Overflow.

### v2

@@ 3 Results
3.1 Validation in job ads
How well do these software skills capture real-world development work beyond SO? To find out, we examine all 50,998 job ads posted to the Hacker News website between January 2018 and May 2024. These ads are unstructured job descriptions, sometimes including salary and expected work activities (tasks). We prompt an LLM (ChatGPT-3.5) to extract salary information and work activities (see Appendix C.1), which yields about 46,000 jobs listing at least 3 work activities, and 3,949 jobs that contain at least 1 work activity as well as salary information. When ads describe more than one position, the LLM provides information on each job separately.
Next, we match each extracted work activity to the closest skill in our taxonomy. To do so, we embed work activities and skills in a 384-dimensional vector space using sBERT (Reimers and Gurevych, 2019). For work activities, we embed the phrases extracted by the LLM; for skills, we calculate the embedding of their labels. [^4] To increase accuracy, we use somewhat longer labels of up to 30 words generated by ChatGPT-4.0. Finally, we match each work activity in a job ad to the closest skill based on the cosine similarity between the embeddings of the work activity and the skills. When we cannot find any skill with a cosine similarity of at least 0.3, we drop the work activity. This yields a 237-dimensional vector of zeros and ones that describe each job’s requirements in terms of the skills of our taxonomy. [^5] In Appendix C.3, we show that our findings are robustness when we change this threshold or when use embeddings for skills based directly on their tags instead of their labels. We combine this vector with information on skill-values and relatedness among skills. Fig. 2a illustrates this process.
@@ 3.1 Validation in job ads
We use these data in two ways to test the validity of our skill taxonomy. Our first validation focuses on the set of advertised jobs that also post salary information. It estimates the association between the average value of a job’s skills and the posted wage in the job ad. We find that skill requirements help predict wage offers (Fig. 2b). Salaries in job ads increase alongside the average value of the skills a job requires. Regression analysis confirms this: a 10% rise in the average value of these skills is associated with a 9% increase in the wage offer for the advertised job (see Appendix C.2).
Second, we ask whether we can predict the skill composition of a job. To do so, we take the skill vectors associated with each job ad and stack these vectors into a dataset that contains for each combination of a skill and an advertised job whether or not the skill is required in the job. Next, we mask 40% of observations in this dataset and calculate the average relatedness of a masked skill requirement, $\theta$, to all unmasked skill requirements in the same job, $j$:
$\mathrm{D}_{\theta,j}=\bm{\bar{\mathrm{r}}}_{\theta}^{j}\bm{\mathrm{s}}^{j},$ (5)
where $\bm{\mathrm{s}}^{j}$ is a vector that subsets the unmasked (i.e., “visible”) portions of job $j$’s skill requirement vector, $S_{j}$. $\bm{\bar{\mathrm{r}}}_{\theta}^{j}$ refers to the corresponding row-normalized vector of relatedness matrix $\mathrm{R}$, consisting of elements $\frac{\mathrm{R}_{\theta,\zeta}}{\sum_{\tau\in\Xi_{j}}\mathrm{R}_{\theta,\tau}}$, where $\Xi_{j}$, the set of unmasked skills in the skill requirements vector of job $j$, such that $\zeta\in\Xi_{j}$ and $\theta\notin\Xi_{j}$.
The skill space strongly predicts which skill requirements appear together in job ads. To show this, Fig. 2c plots how the estimated probability that a job requires a particular (masked) skill changes with that skill’s density in the job’s unmasked skill requirements. First, we sort all masked observations by the unmasked skill’s density in the job, then bin them in 10 equally sized bins and calculate relative frequencies. Next, we plot these estimated probabilities on the vertical axis against the bin’s average skill density on the horizontal axis. 95% confidence intervals are approximated with a normal distribution, using the following estimate of the standard error of the mean:
$\hat{\sigma}_{b}=\frac{1}{\sqrt{n_{b}}}\sqrt{\hat{\pi}_{b}(1-\hat{\pi}_{b})},$ (6)
where $n_{b}$ the number of observations in bin $b$ and $\hat{\pi}=\frac{d_{b}}{n_{b}}$ (where $d_{b}$ the number of skills required in jobs in bin $b$) representing the relative frequency with which masked skills are required in bin $b$.
The probability that a masked skill is required (i.e., has a “1” in the skill requirements vector) increases sharply with its relatedness to the unmasked skill requirements, rising almost 20-fold from 1.0% for unrelated skills to 19.6% for strongly related ones. This not only shows that our SO-based skill taxonomy is meaningful in the real-world software development labor market, but also that software development jobs are coherent: they, on average, require skill combinations that are often mastered by the same users on SO.
Figure 2: Job ads. a. Schematic representation of the workflow to extract salary and work activities from online job ads by prompting ChatGPT. Work activities are converted to the 237-dimensional skill requirement vectors based on cosine similarities between text embeddings of work activities and skills. b. Prediction of wage offers. Jobs are grouped into equally sized bins based on the average value of required skills. The vertical axis shows the estimated mean of the advertised wage offers. c. Prediction of skill requirements. 40% of elements in each skill requirement vector are masked and grouped into 10 equally sized bins based on the average proximity of a masked skill requirement to all unmasked skill requirements in a job. The plot displays the estimated probability that a specific masked skill in a bin is required for a job. Vertical bars in panels b and c represent 95% confidence intervals.
3.2 Mapping skills to programming languages
@@
Programmers can draw on a wide variety of programming languages, each with different levels of suitability and popularity for specific use cases. However, which languages best serve which areas of software development remains hotly debated. Our skill constructs promise a novel way to approach this question: using tags that refer to specific programming languages we can associate SO posts with the languages they refer to (see section 2.2). This, in turn, allows us to track which skills are used with which languages.
@@
Fig. 3a presents a matrix illustrating which programming languages are used by at least 10 users with each of the 237 skills we identified, based on answer postings whose tags refer to both the skill and the language. A visual inspection suggests that the matrix is triangular. This indicates that skills are nested (Hosseinioun et al., 2025), as confirmed by an NODF value of 81.5 (Payrató-Borràs et al., 2020). Accordingly, skills can be roughly ranked from general to specialized based on how widely they are implemented across languages. That is, although different languages may be developed with specific use cases in mind, in practice, this has not led to a strict specialization where dedicated programming languages serve specific areas of software development expertise. Instead, whereas common skills can be performed in almost any language, highly specialized skills often require generalist languages.
@@
Figure 3: Programming languages. a. Skill-language matrix. Elements are colored when at least 10 SO users have at least one answer post in the skill-language combination. b. Number of skills for which a programming language ranks as the top language in terms of SO users. The graph shows time-series for the six largest languages in terms of cumulative SO posts between August 2008 and June 2023. c. Python’s skill footprint. Nodes are colored if Python ranks among the top 3 programming languages for the associated skill. Fully colored: 1st rank, 50% transparency: 2nd rank, 75% transparency: 3rd rank.
@@
Fig. 3b illustrates how different programming languages compete for user types over time. Focusing on the six largest languages on Stack Overflow, it tracks how many distinct skill areas each language dominates by user counts. The most prominent trend is Python’s steady climb: from just nine skills in 2009, it rises to become the top language for over 80 of the 237 skills by 2022, surpassing well-established options like C# and JavaScript. In the Appendix F, we further demonstrate how languages compete for dominance in expertise areas by focusing on the abrupt shift in iOS programming from Objective-C to Swift.
@@
Fig. 3c visualizes Python’s growth, highlighting all skills for which it ranks among the top three languages. Initially focused on data science, Python spread between 2014 and 2018 into software development areas that range from app development to web-related work, DevOps, and cloud computing. Today, it is the dominant language for at least one skill in every cluster. As detailed in the Appendix E, this pattern is preserved when we re-weight languages by their GitHub script counts — a proxy for real-world language usage — indicating that Python’s reach extends beyond Stack Overflow.
4 User expertise
We can study the expertise of SO users, by analyzing the 237-dimensional skill vectors of eq. (1), which count how many answers a user provides to questions associated with each skill in our taxonomy. These vectors therewith capture the intensity of engagement in different areas of software work.
4.1 Voting
To study whether engagement is related to users’ proficiency in a skill, we analyze how many votes their answers receive. Votes are social feedback from other users indicating appreciation for a post. However, before deciding to post an answer, users will first look at existing answers to see if they can improve or add to them. As a consequence, the sample of answers that users post is highly selected. To avoid that this introduces sample selection bias, we limit ourselves to the first recorded answers for each question. Furthermore, users will post answers to question for which they perceive themselves to be sufficiently knowledgeable. We address this type of sample selection bias in Appendix D.2, where we exploit exogenous variation in activity patterns throughout the 24-hour daily cycle.
To assess the effect on the upvotes a user’s answers receive, we estimate the following regression model:
where $X_{\theta,u(a)}$ is user $u(a)$’s expertise in terms of the number of answers provided to questions involving skill $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided. The next two variables serve as control variables for the general amount of interest a question generates on SO: the first, $A_{q(a)}$, is the number of answers provided to question $q(a)$. The second term counts the number of upvotes across all answers to the question, where $Q_{a}$ is the set of answers provided to question $q(a)$ and $v_{i}$ the number of upvotes that answer $i$ received. $\eta_{a}$ is a disturbance term. Finally, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is a question’s top answer or a continuous variable capturing an answer’s popularity: $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues. In some cases, an answer is associated with multiple skills. Whenever this happens, we duplicate the observation accordingly, such that this answer generates one observation for each skill involved.
We expect that users are particularly apt at answering questions that strongly connect to their prior expertise. Fig. 4a supports this conjecture, showing that there is a statistically significant, positive link between users’ skill expertise and the popularity of their answers. The estimates plotted in the figure indicate that a 10% rise in skill expertise boosts the likelihood of an answer becoming the top-voted solution by 0.1 pp ($\pm 0.004$) and increases the number of votes it receives by 0.14% ($\pm 0.006\%$). In the Appendix D.2, we show that correcting for sample-selection issues increases estimated effects to 0.6 pp ($\pm 0.35$) and 0.9% ($\pm 0.41\%$), respectively. Together, these results corroborate the value of skill vectors as summaries of a user’s expertise.
4.2 Skill acquisition
To study how user learn new skills, we randomly divide the population of users into two equally sized samples. We use the first sample, $U_{1}$, to construct a skill relatedness matrix from co-occurrence patterns of skills in users. The second sample, $U_{2}$, is used to estimate how relatedness affects user activity on SO. To do so, we collect information from all answer posts by these users within rolling two-year intervals. That is, we construct expertise vectors $\vec{S}_{u}^{t}$, where $t$ denotes an interval that stretches across the two calendar years preceding year $t$. For instance, $\vec{S}_{u}^{2018}$ refers to the answers user $u$ provides between January 1, 2016 and December 31, 2017. We collect these vectors year-by-year across all users in sample $U_{2}$ in matrices $T^{t}$.
Analogously to eq. (5), we calculate for all users, $u\in U_{2}$, and years, $t$, their expertise density around each skill, collecting results in user-skill matrix $D^{t}$:
$\mathrm{D}^{t}=\bar{\mathrm{R}}^{U_{1}}\mathrm{T}^{t},$ (8)
where $\bar{\mathrm{R}}^{U_{1}}$ is the row-normalized relatedness matrix, constructed from the skill vectors of users in sample $U_{1}$.
Next, we analyze how SO users develop new expertise over time. To do so, we first test whether the relatedness among skills—reflected in the skill space—predicts which new skills a user will acquire. We take all skills in which user $u$ showed no activity in the two calendar years preceding $t$. These skills represent the set of skills that user $u$ has not yet acquired. We tag all such skills with a $1$ whenever the user posts at least one answer related the the skill in year $t$, and with 0 otherwise. Next, we divide all user-skill combinations where a skill is at risk of being acquired by the user into 10 equally sized bins based on their density values. We then estimate the probability that users acquire a skill in year $t$ as the relative frequency of 1s in each bin. Standard errors are calculated analogously to eq. (6).
Fig. 4b shows how these estimated probabilities rise with the user’s expertise density around a skill. Users are most likely to pick up skills closely related to those they already master, with the probability of acquiring a new skill rising from 0.9% (when the skill is unrelated) to 14.3% (when it is highly related).
Apart from a close alignment with their existing skills, higher rewards may also matter. We therefore next regress a dichotomous variable that encodes skill acquisition on the skill value estimated in eq. (3) and the user’s density around the skill:
$y_{u,\theta,t}=\beta_{V}\log(V_{\theta})+\beta_{D}D_{u,\theta,t}+\sigma_{u,\theta,t},$ (9)
where $y_{u,\theta,t}$ is a binary variable that encodes whether or not user $u$ acquires skill $\theta$ in year $t$, $\log(V_{\theta})$ the logarithm of the skill’s value, and $D_{u,\theta,t}$ the user $u$’s density around skill $\theta$ in the two years preceding year $t$, as defined in eq. (8). We estimate this model once with data from all programming languages and once with data that rely only on answer posts related to Python.
Figure 4: a. Regression analysis of answer popularity on skill expertise. Popularity is measured as the number of votes the answer receives, $\log(\#\text{votes}_{a}+1)$ or whether or not the answer is the top answer to the question, skill expertise as the number of prior answers the user provided to questions requiring skill $\theta$ in the preceding two years, $\log(\#\text{answers}_{u(a),\theta,t(a)})$. Control variables include $\log(\#\text{answers}_{q(a)})$ and $\log\left(\sum_{\alpha\in A_{q}}\#\text{votes}_{\alpha}+1\right)$, the total number of answers provided to question $q(a)$, and the sum of all votes across these answers. To avoid problems due to $\log 0$ values, we add 1 to counts that can evaluate to 0. The plot shows point estimates with their 95% confidence intervals. Orange markers refer regression analyses of whether or not an answer is the top answer, purple markers to analyses of the number of votes. b. Estimated probability of acquiring new skills at different values of density, users’ relatedness-weighted expertise in other skills: $d_{\theta,u}=\sum_{\kappa}\frac{\mathrm{R}_{\theta,\kappa}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}X_{\kappa,u}$, where $X_{\kappa,u}$ denotes user $u$’s prior expertise in skill $\kappa$ and $\mathrm{R}_{\theta,\kappa}$ the relatedness between skills $\theta$ and $\kappa$. c. Effect of skill value on the probability that a user acquires a skill from a regression analysis of whether or not a user will acquire a new skill $\theta$, on the skill’s value, $V_{\theta}$, controlling for the user’s density around the skill, $d_{\theta,u}$. Orange markers refer to analyses that use Python-related questions only, purple markers to the full sample.
Findings are summarized in Fig. 4c, with detailed results described in Appendix D.3. Somewhat surprisingly, developers do not universally pursue high-value skills. As shown in the figure’s right-most panel, this result holds even when we control for how related these skills are to the user’s current skill sets. This may reflect higher technological barriers or learning costs associated with high-value skills. Interestingly, however, Python users follow a different pattern: when we restrict the analysis to Python-related questions (shown by purple markers), skill value has a positive effect on their decision to learn a new skill. This suggests that Python’s versatility allows its users to more readily acquire profitable expertise. This not only offers an explanation for Python’s popularity, but also highlights its distinct role in the software ecosystem: as a general-purpose language, it enables users to branch into new, higher-value expertise areas more easily, altering and expanding their career trajectories.